# floZ: Evidence estimation from posterior samples with normalizing flows

### v1

@@ Abstract
We propose a novel method (floZ), based on normalizing flows, for estimating the Bayesian evidence (and its numerical uncertainty) from a set of samples drawn from the unnormalized posterior distribution. We validate it on distributions whose evidence is known analytically, up to 15 parameter space dimensions, and compare with two state-of-the-art techniques for estimating the evidence: nested sampling (which computes the evidence as its main target) and a $k$-nearest-neighbors technique that produces evidence estimates from posterior samples. Provided representative samples from the target posterior are available, our method is more robust to posterior distributions with sharp features, especially in higher dimensions. It has wide applicability, e.g., to estimate the evidence from variational inference, Markov-chain Monte Carlo samples, or any other method that delivers samples from the unnormalized posterior density.
@@ I Introduction
One of the most important scientific tasks is that of discriminating between competing explanatory hypotheses for the data at hand. In classical statistics, this is accomplished by means of a hypothesis test, in which a null hypothesis (e.g., that the data do not contain a signal of interest) is rejected if, under the null, the probability of observing data as extreme or more extreme than what has been observed is small (the so-called $p$-value). An alternative –and often more fruitful– viewpoint is offered by a Bayesian approach to probability, in which the focus is shifted away from rejecting a null hypothesis to comparing alternative explanations (see 2008ConPh..49…71T for a review). This is done by means of the posterior odds (ratio of probabilities) between two (or several) competing models. The central quantity for this calculation is the Bayesian evidence, which gives the marginal likelihood for the data under each model once all the model parameters have been integrated out, tensioning each model’s quality of fit against a quantitative notion of “Occam’s razor”. Bayesian model comparison has received great attention in cosmology (e.g. Martin:2013nzq ; Heavens:2018adv ; 2017PhRvL.119j1301H ), and is being adopted as a standard approach to model selection in the gravitational waves community (e.g. DelPozzo:2011pg ; Vigeland:2013zwa ; Toubiana:2021iuw ) as well as in the exoplanets one (e.g. Lavie_2017 ; 2022ApJ…930..136L ).
@@
Other methods to compute the evidence range from the analytical (Laplace approximation and higher-order moments 10.1214/16-EJS1218 , Savage-Dickey density ratio Trotta:2005ar ) to the ones based on variations of density estimation, like parallel tempering MCMC coupled with thermodynamic integration 10.1063/1.1751356 , and, more recently, on simulation-based inference with neural density (or neural ratio) estimation and/or deep learning 2023arXiv231115650K ; Jeffrey:2023stk ; 2020arXiv200410629R . A promising approach is that of evaluating the Bayesian evidence integral from a set of existing posterior samples, previously gathered e.g. via MCMC. The interest lies in the ability of obtaining the evidence from post-processing of posterior samples, which can be obtained with any suitable algorithm. Such a method, based on the distribution statistics of nearest neighbors of posterior samples, was presented in Heavens:2017afc (see also 2021arXiv211112720M for a similar approach based on harmonic reweighting of posterior samples).
Here, we propose a new method based on normalizing flows. Normalizing flows are a type of generative learning algorithms that aim to define a bijective transformation of a simple probability distribution into a more complex distribution by a sequence of invertible and differentiable mappings. Normalizing flows have been initially introduced in Ref. Tabak:2010 ; Tabak:2013 and then extended in various works, with applications to clustering classification Agnelli:2010 , density estimation Rippel:2013 ; Laurence:2014 , and variational inference Rezende:2015 . While normalizing flows have been used in several parameter estimation scenarios in cosmology and gravitational wave physics Green:2020hst ; Green:2020dnx ; Dax:2021tsq ; Dax:2021myb ; 2023PhRvL.130q1403D ; Wildberger:2022agw ; Bhardwaj:2023xph ; Crisostomi:2023tle ; Leyde:2023iof ; Shih:2023jme , here we introduce them for the first time as a method to evaluate the evidence. This paper is structured as follows: we introduce normalizing flows in Section II, and how a suitable loss can be defined in order to encode the objective of evidence estimation; in Section III, we validate our approach on tractable likelihood in up to 15 parameter space dimensions, and benchmark it against dynesty dynesty_code (an implementation of nested sampling) and the $k$-nearest neighbors method of Heavens:2017afc , demonstrating the superiority of our approach over the latter. We conclude in Section IV with an outlook on future applications.
@@ II.3 Additional loss terms
In our implementation, we exploit all the loss functions described above, following a scheme that we present in the next section (Sec. II.4). This provides better results than using any of the loss terms uniquely or in combination.
@@ III.1 Posterior samples with tractable evidence
  • Truncated $d$-dimensional mixture of five Gaussian distributions: The unnormalized posterior is a mixture model of five different multivariate Gaussians, with equal mixture weights. Each of the five Gaussians has a different mean and covariance matrix, and the distribution is truncated by the rectangular prior. The resulting distribution is non-trivial to model, as discussed in Sec. III.3.
@@ III.2 Posterior samples with intractable evidence
We also test the method with an unnormalized probability distribution with analytically intractable evidence for dimensions higher than 2. For dimensions higher than 3 the evidence is very expensive to compute even numerically, therefore we do not provide a ground truth as reference.
@@
Figure 1 shows examples of the three distributions in 2-dimensions. The Gaussian distribution has mean
@@
respectively. The Rosenbrock case has A = 100, B = 20.
@@
Table 1: Summary of the unnormalized posteriors (and their integrals) used to validate and benchmark our evidence estimation. The lower and upper bounds of the rectangular prior are given by $\bm{{x}}_{i},\bm{x}_{f}$, respectively, the indicator function $I(\bm{z})=1$ for $\bm{z}>\bm{0}$ and 0 elsewhere and $\text{erf}({\bm{x}},\bm{\sigma})$ is the multivariate Normal error function. The vectors $\bm{\mu},\bm{\mu}_{j}\,(j=1,\dots,5)$, covariance matrices $\bm{\sigma},\bm{\sigma}_{j}\,(j=1,\dots,5)\in\mathbb{R}^{d\times d}$ and real scalars $A,B$ are fixed.
@@ III.3 Benchmarking
We evaluate the accuracy and efficiency of our flow-based evidence estimation in comparison to the importance nested sampling technique discussed in Ref. dynesty , using the code in Ref. dynesty_code , and to a $k$-nearest neighbors (hereon, referred to as $k$NN) technique Heavens:2017afc to estimate the evidence from MCMC chains of posterior samples using the recommended $k$ = 1 value. In the three panels of Figure 2, we show the comparison for the 2-dimensional unnormalized probability functions, with a flat rectangular prior. We use $10^{4}$ samples to evaluate the evidence using our method and the $k$NN technique. The nested sampling technique generates about $3\cdot 10^{4}$ samples for the evaluation. For $d=10$ and $d=15$, we evaluate the evidence using $10^{5}$ samples with our method and the $k$NN technique, whereas nested sampling uses between 5 to $9\cdot 10^{4}$ samples. We use the default values of the nested sampler’s tolerance and live points, which are 0.01 and 500 respectively.
Figure 2: Evidence estimation in $d=2$ dimensions for (from left): multivariate Gaussian; finite multivariate Gaussian mixture; Rosenbrock. In all cases, floZ and $k$NN employ $10^{4}$ posterior samples. The true value, represented by the dashed line, has been rescaled to 0, and the shaded regions represent the 1-$\sigma$ uncertainty.
@@
Figure 4: The relative error in the log evidence estimation for different distributions in each panel as a function of the dimensionality of parameter space, $d$. We compare the results of floZ and $k$NN using the same MCMC samples. We also show the evidence from nested sampling (labeled as NS). Since the ground truth for the higher-dimensional Rosenbrock likelihoods (grey-shaded region) is not numerically tractable, we compare the relative deviation from the floZ mean prediction (shown by the secondary y-axis).
Figure 4 compares the evidence evaluated by floZ, $k$NN, and nested sampling (labeled NS) techniques for the three distributions, each simulated in 2, 10, and 15 dimensions. The single Gaussian scenario has the simplest distribution and as a result, all methods perform well. However, with increasing complexity and higher dimensions, floZ and nested sampling are both more accurate than $k$NN.
@@
It is difficult to fairly compare the computational effort required by floZ versus nested sampling, as the former takes as input posterior samples previously generated. Moreover, the efficiency of nested sampling depends on the details of the algorithm used, as well as on its settings (typically, the number of live samples and tolerance adopted). While we leave a more detailed study of actual scientific cases to future work, the $d=15$ results of this paper were obtained in less than 10 minutes on an A100 GPU.
IV Conclusions

### v2

# floZ: Improved Bayesian evidence estimation from posterior samples with normalizing flows

@@ Abstract
We introduce floZ, an improved method based on normalizing flows, for estimating the Bayesian evidence (and its numerical uncertainty) from a set of samples drawn from the unnormalized posterior distribution. We validate it on distributions whose evidence is known analytically, up to 15 parameter space dimensions, and compare with two state-of-the-art techniques for estimating the evidence: nested sampling (which computes the evidence as its main target) and a $k$-nearest-neighbors technique that produces evidence estimates from posterior samples. Provided representative samples from the target posterior are available, our method is more robust to posterior distributions with sharp features, especially in higher dimensions. For a simple multivariate Gaussian, we demonstrate its accuracy for up to 200 dimensions with $10^{5}$ posterior samples. floZ has wide applicability, e.g., to estimate the evidence from variational inference, Markov Chain Monte Carlo samples, or any other method that delivers samples from the unnormalized posterior density, such as simulation-based inference. We apply floZ to compute the Bayes factor for the presence of the first overtone in the ringdown signal of the gravitational wave data of GW150914, finding good agreement with nested sampling.
@@ I Introduction
One of the most important scientific tasks is that of discriminating between competing explanatory hypotheses for the data at hand. In classical statistics, this is accomplished by means of a hypothesis test, in which a null hypothesis (e.g., that the data do not contain a signal of interest) is rejected if, under the null, the sampling probability of obtaining a summary statistics as extreme or more extreme than what has been observed is small (the so-called $p$-value). An alternative –and often more fruitful– viewpoint is offered by a Bayesian approach to probability, in which the focus is shifted away from rejecting a null hypothesis to comparing alternative explanations (see 2008ConPh..49…71T for a review). This is done by means of the posterior odds (ratio of probabilities) between two (or several) competing models. The central quantity for this calculation is the Bayesian evidence, which gives the marginal likelihood for the data under each model once all the model parameters have been integrated out, tensioning each model’s quality of fit against a quantitative notion of “Occam’s razor”. Bayesian model comparison has received great attention in cosmology (e.g. Martin:2013nzq ; Heavens:2018adv ; 2017PhRvL.119j1301H ), and is being adopted as a standard approach to model selection in the gravitational wave (GW) community (e.g. DelPozzo:2011pg ; Vigeland:2013zwa ; Toubiana:2021iuw ; Roulet:2024cvl ) as well as in the exoplanets one (e.g. Lavie_2017 ; 2022ApJ…930..136L ).
@@
Other methods to compute the evidence range from the analytical (Laplace approximation and higher-order moments 10.1214/16-EJS1218 , Savage-Dickey density ratio Trotta:2005ar ) to the ones based on variations of density estimation, like parallel tempering MCMC coupled with thermodynamic integration 10.1063/1.1751356 and, more recently, on simulation-based inference with neural density (or neural ratio) estimation and/or deep learning 2023arXiv231115650K ; Jeffrey:2023stk ; 2020arXiv200410629R . A promising approach is that of evaluating the Bayesian evidence integral from a set of existing posterior samples, previously gathered e.g. via MCMC. The interest lies in the ability to obtain the evidence from the post-processing of posterior samples, which can be obtained with any suitable algorithm. Such a method, based on the distribution statistics of nearest neighbors of posterior samples, was presented in Heavens:2017afc (see also Ref. 2021arXiv211112720M for a similar approach based on harmonic reweighting of posterior samples, recently upgraded to a normalizing-flows based method in Piras2024TheFO , and Ref. Rinaldi:2024bow for a hierarchical inference approach).
As first pointed out by JiaSeljak_NFEvidence_2019 , the evidence can also be extracted using normalizing flows. Building on this idea, we present an improved algorithm that provides a fast, accurate, and robust estimate of the evidence from posterior samples. Normalizing flows are a type of generative learning algorithms that aim to define a bijective transformation of a simple probability distribution into a more complex distribution by a sequence of invertible and differentiable mappings. Normalizing flows have been initially introduced in Ref. Tabak:2010 ; Tabak:2013 and then extended in various works, with applications to clustering classification Agnelli:2010 , density estimation Rippel:2013 ; Laurence:2014 , and variational inference Rezende:2015 . While normalizing flows have been used in several parameter estimation scenarios in cosmology and GW physics Green:2020hst ; Green:2020dnx ; Dax:2021tsq ; Dax:2021myb ; 2023PhRvL.130q1403D ; Wildberger:2022agw ; Bhardwaj:2023xph ; Crisostomi:2023tle ; Leyde:2023iof ; Shih:2023jme ; Vallisneri:2024xfk , here we introduce a novel training prescription that is specifically tailored towards evaluating the evidence. This paper is structured as follows: we introduce normalizing flows in Section II, and how a suitable loss can be defined in order to encode the objective of evidence estimation; in Section III, we validate our approach on tractable likelihood in up to 15 parameter space dimensions, and benchmark it against dynesty dynesty_code (an implementation of nested sampling) and the $k$-nearest neighbors method of Heavens:2017afc , demonstrating the superiority of our approach over the latter. In Section IV, we apply our method to real GW data analysis by estimating the Bayes factor in favor of the presence of the first overtone in the ringdown phase of GW150914. We conclude in Section V with an outlook on future applications.
@@ II.3 Additional loss terms
In our implementation, we exploit all the loss functions described above, following a scheme that we present in the next section (Sec. II.4). This provides higher accuracy than using any of the loss terms uniquely or in combination.
@@ II.4 Algorithm
A perfectly trained flow transforms the real samples into a normal distribution with zero mean and unit variance in latent space, which in future work we will exploit to define a stopping criterion for training. This could be achieved by verifying that the flow’s latent distribution is congruent to a standard Normal, for example, via the Kullback–Leibler divergence.
II.5 Addressing limitations of normalizing flows
Figure 1: Effect of high target density near prior boundaries on the normalizing flow prediction. The color plot compares the original, “sharp” boundary distribution (left) with the reflected distribution (right). The color bar represents the un-normalized posterior probability density. The grey lines highlight the boundary. The background black (left) and green (right) scatter points are the predicted distribution by the flow trained on the respective distributions.
Although normalizing flows are very expressive in modeling complex distributions, in the context of evidence estimation we find they can be inefficient in two aspects.
  • Modeling distributions that lie far from the origin: As the latent space is centered at the origin, the flow can require many training cycles to correct for the translation required to model the target real distribution. We reduce the training time by subtracting the mean value from the samples, thus re-centering the bulk of the real distribution to the origin.
  • Modeling distributions with high density in proximity to boundaries: Flows model a target distribution through a diffeomorphic transformation to the latent space. Therefore, they inherently struggle to model discontinuities, i.e. distributions with a high density near a sharp prior boundary. We circumvent this limitation by sequentially transferring half of the samples to their reflected position about each of the sharp edges, ultimately resulting in a continuous distribution better suited for the flow to model. About each sharp edge, the prior volume is doubled, thus halving the prior density. To account for the change, we rescale the un-normalized probability of all samples by a factor of 2 for each sharp edge.
    Figure 1, left, shows one such distribution with peak density adjacent to two boundaries, marked by the grey lines. We perform successive sample transfers about both boundaries, resulting in the distribution on the right. Note that the un-normalized probability of the final distribution is one-fourth that of the original due to the quadrupling of the prior volume. The discrepancy between the target distribution (color plot) and the flow’s predicted distribution (scatter points) is apparent in the original distribution, as illustrated by the leakage of the flow’s predicted samples outside the prior boundaries. In contrast, the reflected distribution has flow samples consistent with the target, with no significant outliers.
    Figure 2: Effect of sharp boundaries on the evidence estimation. The black (green) probability density shows the distribution of the log evidence prediction from the original (reflected) samples. The log evidence is re-scaled by the ground truth, so the true value is at 0 (dashed black line).
    Figure 2 compares the evidence estimation from the two distributions. The prediction from the original distribution significantly overestimates the evidence, while the estimation using reflected samples is well-centered on the ground truth.
    Our method of translating samples about sharp edges can be adapted for other techniques that estimate the evidence from samples, such as techniques that use nested sampling and k-nearest neighbor. We discuss this with an example in Sec III.3 and Figure 6.
    Finally, in the case of periodic parameters (such as angles), a distribution may have an over-density at the upper and lower boundary. We smoothly merge these densities by appropriately translating the parameter samples, modulo the periodicity.
@@ III.1 Posterior samples with tractable evidence
  • Truncated $d$-dimensional mixture of five Gaussian distributions: The unnormalized posterior is a mixture model of five different multivariate Gaussians, with equal mixture weights. Each of the five Gaussians has a different mean and covariance matrix, and the distribution is truncated by the rectangular prior. The Gaussian mixture is a non-trivial extension of the single Gaussian case.
  • Truncated $d$-dimensional exponential function: We define a multivariate, unnormalized posterior that follows an exponential shape along positive values of each parameter, and enforce a rectangular prior with one vertex at the origin. We assign random (but fixed) values to the components of the exponent $\bm{\lambda}$. The unnormalized posterior is peaked at the origin. The resulting sharp distribution is non-trivial to model, as discussed in Sec. II.5. We further highlight the importance of addressing sharp posterior boundaries in Sec. III.3.
@@ III.2 Posterior samples with intractable evidence
We also test the method with an unnormalized probability distribution with analytically intractable evidence for dimensions higher than 2. For dimensions higher than 3 the evidence is very expensive to compute even numerically, therefore we do not provide a ground truth as a reference.
@@
Figure 3 shows examples of the four distributions in 2-dimensions. The Gaussian distribution has mean
@@
respectively. The exponential distribution has
$\bm{\lambda}=\begin{pmatrix}0.009057\\ 0.005257\end{pmatrix}.$
The Rosenbrock case has A = 100, B = 20.
@@
Exponential $\mathrm{exp}\left[-\bm{\lambda}\,\bm{x}\right]I({\bm{x}})I({\bm{x}}_{f}-{\bm{x}})$ $|\prod_{i=1}^{n}\lambda_{i}|^{-1}\left[1-\mathrm{exp}(-\bm{\lambda}\,\bm{x}_{f})\right]$
@@
Table 1: Summary of the unnormalized posteriors (and their integrals) used to validate and benchmark our evidence estimation. The lower and upper bounds of the rectangular prior are given by $\bm{{x}}_{i},\bm{x}_{f}$, respectively, the indicator function $I(\bm{z})=1$ for $\bm{z}>\bm{0}$ and 0 elsewhere and $\text{erf}({\bm{x}},\bm{\sigma})$ is the multivariate Normal error function. The vectors $\bm{\mu},\bm{\mu}_{j}\,(j=1,\dots,5)$, $\bm{\lambda}$, covariance matrices $\bm{\sigma},\bm{\sigma}_{j}\,(j=1,\dots,5)\in\mathbb{R}^{d\times d}$, and real scalars $A,B$ are fixed.
@@ III.3 Benchmarking
Figure 4: Evidence estimation in $d=2$ dimensions for (clockwise from top-left): multivariate Gaussian; finite multivariate Gaussian mixture; Rosenbrock; Exponential. In all cases, floZ and $k$NN employ $10^{4}$ posterior samples. The true value, represented by the dashed line, has been rescaled to 0, and the shaded regions represent the 1-$\sigma$ uncertainty.
@@
Figure 6: The relative error in the log evidence estimation for different distributions in each panel as a function of the dimensionality of parameter space, $d$. We compare the results of floZ and $k$NN using the same MCMC samples. We also show the evidence from nested sampling (labeled NS). For the exponential case, the results with solid shading were computed after processing the samples around the sharp edges, whereas those with translucent shading were computed with the sharp distribution edges (labeled ’Sharp’). Since the ground truth for the higher-dimensional Rosenbrock likelihoods (grey-shaded region) is not numerically tractable, we compare the relative deviation from the floZ mean prediction (shown by the secondary y-axis).
Figure 7: High dimensional scalability: the relative error in the log evidence estimation for the single Gaussian case as a function of the dimensionality of parameter space, $d$. For all dimensions, we use $10^{5}$ posterior samples. We use 11 hidden layers for $d=10$ and 20 hidden layers for higher dimensions.
@@
Figure 6 compares the evidence evaluated by floZ, $k$NN, and nested sampling (labeled NS) techniques for the four distributions, each simulated in 2, 10, and 15 dimensions. For $d=10$ and $d=15$, we evaluate the evidence using $10^{5}$ samples with our method and the $k$NN technique, whereas nested sampling uses between 5 to $9\cdot 10^{4}$ samples. For the exponential case, due to its sharp distribution. we show the results with (solid colors) and without (translucent colors) our prescribed method of sample translation about sharp edges (Sec II.5).
The single Gaussian scenario has the simplest distribution and as a result, all methods perform well. However, with increasing complexity and higher dimensions, floZ and nested sampling are both more accurate than $k$NN.
The exponential distribution highlights the challenge that sharp distributions pose to accurate density and evidence estimation, exacerbated by increasing dimensions. All three techniques show marked improvement upon performing the translation. In addition to a more accurate estimation, the error bars of $floZ$ decrease.
@@
It is difficult to compare the computational effort required by floZ versus nested sampling, as the former takes as input posterior samples previously generated. Moreover, the efficiency of nested sampling depends on the details of the algorithm used, as well as on its settings (typically, the number of live samples and tolerance adopted). For what concerns floZ, the $d=15$ results of this paper were obtained in less than 10 minutes on an A100 GPU.
In Figure 7, we demonstrate the scalability of $floZ$ up to 200 dimensions in the case of a single multivariate Gaussian, whose evidence has been estimated with a constant number of posterior samples, $10^{5}$ for all cases. The single Gaussian is of course the easiest target for normalizing flows, as the transformation onto the latent space is trivial; more challenging posteriors are not expected to train as well in many dimensions, and would require a larger number of training examples. However, for mildly non-Gaussian distributions of the kind often considered in cosmological data analysis, we expect that floZ can deliver accurate evidence estimation for parameter space of up to $O(100)$ dimensions.
IV Application to real data: gravitational wave analysis
Figure 8: The relative log evidence estimation for the fundamental mode ($\mathrm{QNM}_{220}$) only, left, and also the first overtone ($\mathrm{QNM}_{220}+\mathrm{QNM}_{221}$), center. The evidence is rescaled by the mean of the nested sampling (NS) estimate. The Bayes factor (B) in favor of the presence of the first overtone is shown on the right. The shaded regions represent the $1\sigma$ uncertainty.
As a practical application of our method in astrophysics, we investigate the presence of the first overtone in the ringdown phase of the first GW detection GW150914 GW150914 . The ringdown signal from a binary black hole merger can be decomposed as a superposition of quasi-normal modes (QNMs), whose complex frequencies are uniquely determined by the final mass and spin of the remnant black hole. Therefore, the observation of more than a single QNM allows for performing consistency tests of General Relativity 1980ApJ…239..292D ; Dreyer:2003bv ; Berti:2005ys .
While it is established that GW150914 contains the fundamental ($n=0$, $l=m=2$) mode 2016PhRvL.116v1101A , there are conflicting claims regarding the presence of the first overtone ($n=1$, $l=m=2$). Ref. 2019PhRvL.123k1102I found the first overtone to be present with a $3.6\sigma$ confidence. On the contrary, Ref. 2022PhRvL.129k1102C found no evidence in favor of an overtone in the data, with a Bayes factor around the waveform peak lower than one. The discussion between the authors of 2019PhRvL.123k1102I and 2022PhRvL.129k1102C is still ongoing Isi:2022mhy ; Isi:2023nif ; Carullo:2023gtf , and the debate is still unsettled. In the meantime, independent analyses (using different techniques) provided different results: Refs. 2022PhRvD.106d3005F ; Crisostomi:2023tle ; Wang:2023mst reported a lower significance for the presence of the overtone (which does not reach $2\sigma$); Ref. Wang:2023xsy found the Bayes factor for the overtone to lie between 10 and 26; Ref. Correia:2023bfn a Bayes factor of 2.3; and Ref. 2023PhRvD.107h4010M a Bayes factor of 600.
These discrepant results can be attributed to the sensitivity of the analysis to different assumptions such as the detector noise characterization, ringdown start time, length of the signal to analyze, marginalizing vs. fixing some of the parameters, and presence of non-linear effects. The aim of this section is neither to address the aforementioned issues, nor to provide yet another claim regarding the evidence of the first overtone in the data (or lack thereof), but we rather aim to test the performance of floZ in a real scientific case of interest.
[…]
pyRing relies on the nested sampling algorithm CPNest cpnest , which also provides the evidence estimation we use to compare to our for validation (for comparison, we also estimate the evidence with the $k$NN technique). We used 4096 live points, 4096 maximum Markov Chain steps, and a tolerance of 0.01. For the one-mode model CPNest provides 17250 samples, instead for the two-mode model 20000 samples. Since the posterior distribution for the phase of the fundamental mode peaks close to the boundaries, following the method of periodic translations described in Sec II.5, we shift this parameter to make the distribution unimodal.
Figure 8 shows the evidence estimated by floZ and $k$NN scaled by the result from CPNest. The Bayes factor in favor of the presence of the first overtone, as obtained by $floZ$, is $0.87\pm 0.16$. For comparison, nested sampling gives $0.76\pm 0.15$ and $k$NN $1.49\pm 0.02$. The estimates obtained via nested sampling and floZ are compatible within their $1\sigma$ uncertainties.
As a final note, let us stress that although in this particular case the use of floZ would seem unnecessary since a nested sampler is already available, the running time of pyRing, on 4 Intel(R) Xeon(R) Gold 6238R CPU @ 2.20GHz, was of the order of several hours, which should be compared to the few minutes required by floZ. Therefore, for ringdown analyses, our tool could be better paired with the latest implementation of ringdown maximiliano_isi_2021_5094068 ; Isi:2021iql on a JAX back-end (which supports GPU execution) and the Hamiltonian Monte Carlo (HMC) sampler in its NumPyro implementation, which could supply posterior samples from the unnormalized target for fast evidende estimation with floZ.
V Conclusions